18. Which of the vectors vi V3 are (a) orthogonal? (b) in the same direction? (c) in opposite directions? -13
Answer #1

18 . Å Vi= 1 î -1% - 2 2 = 2? +49-1 V₂ = 3 i 1 û + 2 ? ê VF i î ro f t I was (a) Orthogonal for being orthogonal å bo o4ia of(6) corr= az a= x + ayt a cos a= ane cosß= ay jai 191 For a vertorb to be in same direction b = 6(9x + ay + O₂) bx= banc, by=